# Motivation & Scientific Context

### v1

@@
Despite this positive growth of projects and initiatives, developing astronomy research in Africa faces many on-going challenges. These are shared across many research topics and institutions across the continent, and for people of African origin (but may be working in other countries). These relate to: limited opportunities and resources; the limited retention of young people into higher education; issues with equity and inclusion (including challenges around travel visas); and access to the latest international knowledge, facilities, and networks [4, 5].
@@
Both observational studies and simulations have shown that the energy released by AGN is important for galaxy evolution. However, there continue to be major research challenges in the field, including: (1) obtaining a complete census of AGN events; (2) understanding the detailed physical structure of the material associated with AGN; (3) establishing the properties of AGN host galaxies and the details of how AGN regulate galaxy growth, and; (4) identifying the key techniques and datasets required to make significant progress in answering these scientific questions over the coming decade [review in e.g., 6]. The scientific scope of our conference was to cover these four main themes. [^5] We note that these themes cover a number of the STFC’s Science Challenges in frontier physics (A1 - A5). Due to the focus on the African scientific community, the workshop looked to particularly showcase work on this scientific topic that uses SKA pathfinder observatories in addition to other astronomical facilities based on the African continent.
@@
Careful planning went into selecting participants. During selection, our primary objectives were: (1) ensuring a scientifically productive meeting, with participants working on the relevant research problems, balanced across the four key themes; and (2) having significant participation from those of African institutions and/or with African origin (aiming for $\approx$30–50%). We detail the processes of selecting the organisational committees, choosing invited speakers, and the selection of participants.
@@
Our Local Organisation Committee (LOC) was put together primarily from the research teams (i.e., postdoctoral researchers and PhD students) of the workshop co-chairs (i.e., Chris Harrison, based at Newcastle University and Leah Morabito, based at Durham University). Therefore, they had a direct scientific interest in the workshop. The LOC were responsible for the typical activities of organising a conference (e.g., coordinating with participants, helping with administrative tasks, supporting the activities during the week etc.). Within the LOC there were 4 researchers with an African origin (representing four different African countries), although they were based at UK institutions at the time. Their perspective was crucial for helping to identify - and overcome - many of the challenges that African participants would face in travelling to the UK, as well as for developing a programme of activities to ensure participants felt included, supported, and comfortable to contribute. A photograph of the LOC is shown in the top panel of Figure A.1.
The Scientific Organising Committee (SOC) were responsible for designing the scientific programme, choosing invited speakers, and selecting the conference participants. We ensured that we had significant representation of scientists at African institutions as well as representatives from the UK and wider Europe. Of the 9 SOC members, 4 were based at African institutions (from 3 separate countries). Furthermore, the selection of the SOC was made to ensure representation of established experts covering a diverse range of the scientific topics that were the focus of the conference. Consequently, the SOC not only provided crucial scientific input, but additionally were able to help design a participant selection process that was suitable for African-affiliated researchers (especially early-career researchers). Furthermore, they suggested activities for the programme that would be inclusive to those who have little-to-no experience of international conferences. Importantly, these SOC members could locally advertise the conference across the African community. A photograph of the SOC is shown in the bottom panel of Figure A.1.
@@
Along with a dedicated website, we released an application form for conference participation 10 months before the actual conference. The deadline for submitting an application was $\approx$7 months before the conference dates. This long lead time was critical to ensure sufficient time to support travel and visa applications, as well as plan and to conduct pre-conference activities (described later).
@@
We had capacity for 120 participants and an ideal target of $\approx$110. The final number of participants attending the conference was 107, of which 9 could only join online due to not being able to travel after unforeseen circumstances. Figure 1 shows a world map highlighting the countries of origin and affiliation that were represented across the conference participants. Overall, 43% of the participants were either from an African institution and/or of African origin. There were 20 different affiliation countries represented, of which 11 were from Africa. The break down of career stage was: 31% senior academic/lecturer; 26% postdoc/fellow; 31% PhD student; and 12% Masters students. For gender, the final breakdown was 49% Male, 49% Female, and 2% Non-binary. A conference photo, including the majority of participants, is shown in Figure A.2.
@@
Offers for conference participation were sent out $\approx$6 months before the conference itself. This long lead-time was chosen to ensure sufficient time to help participants who needed to apply for visas, and/or whose travel we needed to support. Indeed, visa applications can take a significant amount of time. For example, one participant had a visa issued only two days before travel. For similar reasons, it was necessary to book travel for fully-funded participants before visa applications were completed, otherwise flight costs could have risen to unaffordable rates as the travel dates approached.
@@
The only costs that fully-funded participants were required to pay for themselves up-front, and later obtain a reimbursement, were the visa costs, travel insurance (as often not provided by the home universities), and local travel costs. We note that costs associated with obtaining visas were sometimes quite high, due to participants needing to travel to consulates (sometimes multiple times). Although this process generally went well, some applicants had no access to personal bank accounts and arrangement had to be made with a third party to make the reimbursement because cash payments were not possible. In some cases, we were not aware of it until after it was too late to help with alternative methods of payment beforehand.
@@
The participants applied for and received standard visitor visas. These are valid for 6 months, and some participants organised other work visits while they were in the UK. This required some extra coordination, but we were happy to support it. We had some requests during the meeting or after for other work visits, and it would have been useful to highlight this opportunity to participants beforehand. This could have allowed participants to make the best use of their visas while they were valid. However, we caution that while visits to the UK for conferences and meeting do not require approval from the Academic Technology Approval Scheme (ATAS), research visits of any length do require approval, at the time of writing. If people wish to join a visit to carry out research along with this, we recommend they apply for ATAS certification at the same time or even before the visa.
@@
The following sub-sections describe the activities and their impact, as much as we are able to assess four months after the event.
@@
Around 1 month before the workshop, we created a collaborative online Slack space, to allow participants to start introducing themselves to each other. Additionally, to encourage participants getting to know each other before the conference, we asked participants to create 2-minute introductory videos. These videos were shared on a private YouTube playlist in the month leading up to the conference. Participants were provided with clear guidelines on how to record these videos on a phone, and what these videos should include (i.e., their name, affiliation, research interests, and what they hope to get out of the conference). Only 35/107 participants created these videos; with mostly the more junior participants participating (the LOC and SOC were all compelled to make a video). Nonetheless, we had several positive comments during the feedback questionnaire about the introductory videos being a good idea, and useful to start to get to know fellow participants.
@@
We split into groups of $\approx$25, across separate rooms. These groups were then re-shuffled half way through the session. Before the activity all participants had been assigned their groups/rooms so that they knew where to be. Two activities were used: (1) “Find 10 things in common” and (2) “Queen Bee: Find the Answer to my Questions”. The former activity had each group split into 2-3 sub-groups and then take turns to briefly introduce themselves to each other. They then had 10 minutes to try and find 10 things in common across the whole group. This could be related to research tools used, hobbies, favourite types of food etc. The second activity involved sub-groups of $\approx$3–5. One person in each group is nominated the ‘Queen Bee’, who gives their ‘worker bees’ a question that they want the answer to. This could be something scientific (e.g., ‘who could I talk to about processing these type of astronomical data’), or anything else (e.g., ‘what’s the best restaurant to get good Indian in Newcastle’). All the worker bees, from each group, then spend $\approx$3 minutes speaking to everyone else in the room to get the answers, and report back to the Queen Bee. The Queen Bees could then report back to the whole room what they found out, before somebody else takes a turn as Queen Bee.
The top-left panel of Figure 2 shows the results of asking participants “How did you find the networking session on Monday?”, with answers between 1 for poor and 5 for excellent. It can be seen that the activity was extremely popular across all participants, with 91% scoring 4 or 5. In Figure 2 we also split the responses for only the African affiliated/origin participants, and found equal levels of positive responses for this specific group. The qualitative comments suggested the activity was very effective for helping early-career researchers feel more relaxed and able to approach people for conversations throughout the week. It was especially noticed feeling more comfortable approaching senior academics after this activity. Some other comments suggested the activity was a bit tiring, so it could have been reduced in length and/or more rest time built in to the programme after this intense activity.
@@
Sparkler talks were 90 second presentations (with one slide only). These were done over 2 scheduled blocks on the first two days. The purpose of these were to allow participants to introduce themselves, their main research interest and/or one key scientific result, and to advertise a poster (if relevant). As described above, participants were given examples and training on how to effectively put these together, with a focus on avoiding participants trying to give a full scientific talk very quickly!
The SOC selected chairs for the contribution sessions, to be more senior people, but keeping the geographical representation the same as the overall conference. We provided the chairs with moderator guidelines, which most chairs adhered to. These included information about keeping to time, encouraging a diversity of participants to ask questions, and ensuring the code of conduct is upheld. The LOC provided technical support for each session, with specific LOC members identified for each session.
@@
The feedback we received indicates that we achieved our objective of creating a collaborative environment where people ‘across continents’ could come together to network and share in scientific discussion. The bottom-right panel of Figure 2 shows the results of the participant feedback to the question: “How comfortable did you feel to have discussions/conversations with new people during the conference (e.g., during coffee breaks, lunchtimes…)”, ranging from 5 (Very) to 1 (Not at all). The results are overall positive with 55% scoring 5 and 36% scoring 4, with a similar positive response for only the African affiliated/origin groups (62% scoring 5 and 36% scoring 4).
@@
We received little feedback from the small number of online participants. However, we know that there they mostly only participated during the scientific contributions part of the conference. Our focus was always on generating an effective in-person conference experience. Future consideration would be needed to make a more effective hybrid conference format, and/or to consider an online-only equivalent.
@@
Only four months after the conference, we are not yet able to establish its long-term impact. However, initial signs are encouraging. For example, when we asked “Would you be interested in attending a conference like this in a future?”, (see Figure 3), 94% of all participants answered “yes” (100% of the African affiliated/origin participants). There was enthusiasm to run another similar meeting, potentially on the continent of Africa next time.
@@
We already have informal knowledge of several new collaborations which have resulted from the conference. For example, a Large Programme to study AGN with the Southern African Large Telescope, led by African scientists, including non-African partnerships that were fostered during the workshop. Other smaller examples include two PhD students (one of African origin) deciding to initiate a new project combining the data of one student with the methods of the other student, and a student visit to Durham to start a radio project (combined with a workshop).
We are also aware that many participants who participated in the Buddy Scheme, are still in contact with their buddies, further highlighting the success in forming new networks and working relationships.
We plan to follow up with participants to ask them about collaborations they formed during the conference, in the 6–12 months after the conference. These results will be helpful in providing evidence of any long-term impact that our approach for this conference has had.
@@
The main goal of this conference was to hold a scientific meeting, discussing timely scientific problems, but with a broader impact to integrate African researchers into the international community and to strengthen ties between the European and African scientific communities. According to the feedback from one participant, they reached their scientific goals for the conference, along with the bonus of “1. unplanned research insights and collaboration directions …; 2. appreciating the responsibility that comes with the privilege of being West-affiliated; 3. getting to know African astronomers who are great role models for everyone” From a different perspective, “Having greater representation of Africans really made people much more approachable.” The social activities allowed everyone to bond, while many people noted that “The quality of the participants’ science contributions has been really impressive.” The feedback has been overwhelmingly positive, and there is appetite for another conference. We seem to have struck the balance between science and social aspects about right: as one participant noted the conference was “really well thought out in terms of inclusivity, time table, networking opportunities, making people (especially students) feel comfortable to attend and talk to more senior researchers, super positive and supportive atmosphere, extremely interesting and high-quality science.”
@@
We have learned a lot from this entire process, from simple things like: what do people actually have to pay for if applying for a visa? and what are the logistical considerations we take for granted that may be different in other countries? Even the convention of putting a first name before a family name may be different across countries. The diversity of the LOC and SOC were fundamental in identifying potential issues before they arose, and problem-solving when unexpected challenges came up. The biggest challenges revolved around ensuring our budget would stretch to cover all of the costs we promised, and our key recommendation here is that funding requests start early, and that as much as possible should be paid for directly by the host organisations.
@@
The funding for this conference pulled together from several sources. It started from the fact that the co-chairs had small amounts from their fellowships for modest science meetings, which could be combined for a single, more impactful meeting. Other funding was necessary, and it is crucial to start looking for funding sources early ($\approx$18–24 months before). While we plan on following up the conference participants in future to track long-term collaborations, we would also like to start planning another conference to follow this one. The timeline for this is not yet set, and will likely depend on funding.
@@
Overall, this meeting exceeded our expectations. The LOC and SOC did an incredible job to enact the vision, and to ensure that every small aspect was considered and addressed. Before and during the conference, the engagement from all participants was absolutely amazing. It was wonderful to see senior and junior researchers from Africa, Europe, and beyond, interacting and discussing science. We hope that the collaborations that have started will bring our communities together for a more global approach to science and inclusivity.

### v2

@@
Despite this positive growth of projects and initiatives, developing astronomy research in Africa faces many on-going challenges. These are shared across many research topics and institutions across the continent, and for people of African origin (but may be working in other countries). These relate to: limited opportunities and resources; the limited retention of young people into higher education; issues with equity and inclusion (including challenges around travel visas); and access to the latest international knowledge, facilities and networks [4, 5].
@@
Both observational studies, and simulations, have shown that the energy released by AGN is important for galaxy evolution. However, there continue to be major research challenges in the field, including: (1) obtaining a complete census of AGN events; (2) understanding the detailed physical structure of the material associated with AGN; (3) establishing the properties of AGN host galaxies and the details of how AGN regulate galaxy growth, and; (4) identifying the key techniques and datasets required to make significant progress in answering these scientific questions over the coming decade [review in e.g., 6]. The scientific scope of our conference was to cover these four main themes. [^5] We note that these themes cover a number of the STFC’s Science Challenges in frontier physics (A1 - A5). Due to the focus on the African scientific community, the workshop looked to particularly showcase work on this scientific topic that uses SKA pathfinder observatories in addition to other astronomical facilities based on the African continent.
@@
Careful planning went into selecting participants. During selection, our primary objectives were: (1) ensuring a scientifically productive meeting, with participants working on the relevant research problems, balanced across the four key themes (outlined above); and (2) having significant participation from those of African institutions and/or with African origin (aiming for $\approx$30–50%). We detail the processes of selecting the organisational committees, choosing invited speakers and the selection of participants.
@@
Our Local Organisation Committee (LOC) was put together primarily from the research teams (i.e., postdoctoral researchers and PhD students) of the workshop co-chairs (i.e., Chris Harrison, based at Newcastle University and Leah Morabito, based at Durham University). Therefore, they had a direct scientific interest in the workshop. The LOC were responsible for the typical activities of organising a conference (e.g., coordinating with participants, helping with administrative tasks, supporting the activities during the week etc.). Within the LOC there were four researchers with an African origin (representing four different African countries), although they were based at UK institutions at the time. Their perspective was crucial for helping to identify - and overcome - many of the challenges that African participants would face in travelling to the UK, as well as for developing a programme of activities to ensure participants felt included, supported, and comfortable to contribute. A photograph of the LOC is shown in the top panel of Figure A.1.
The Scientific Organising Committee (SOC) were responsible for designing the scientific programme, choosing invited speakers and selecting the conference participants. We ensured that we had significant representation of scientists at African institutions as well as representatives from the UK and wider Europe. Of the nine SOC members, four were based at African institutions (from three separate countries). Furthermore, the selection of the SOC was made to ensure representation of established experts covering a diverse range of the scientific topics that were the focus of the conference. Consequently, the SOC not only provided crucial scientific input, but additionally were able to help design a participant selection process that was suitable for African-affiliated researchers (especially early-career researchers). Furthermore, they suggested activities for the programme that would be inclusive to those who have little-to-no experience of international conferences. Importantly, these SOC members could locally advertise the conference across the African community. A photograph of the SOC is shown in the bottom panel of Figure A.1.
@@
Along with a dedicated website, we released an application form for conference participation ten months before the actual conference. The deadline for submitting an application was around seven months before the conference dates. This long lead time was critical to ensure sufficient time to support travel and visa applications, as well as plan and to conduct pre-conference activities (described later).
@@
We had capacity for 120 participants and an ideal target of $\approx$110. The final number of participants attending the conference was 107, of which nine could only join online due to not being able to travel after unforeseen circumstances. Figure 1 shows a world map highlighting the countries of origin and affiliation that were represented across the conference participants. Overall, 43% of the participants were either from an African institution and/or of African origin. There were 20 different affiliation countries represented, of which 11 were from Africa. The break down of career stage was: 31% senior academic/lecturer; 26% postdoc/fellow; 31% PhD student; and 12% Masters students. For gender, the final breakdown was 49% Male, 49% Female, and 2% non-binary. A conference photo, including the majority of participants, is shown in Figure A.2.
@@
Offers for conference participation were sent out roughly six months before the conference itself. This long lead-time was chosen to ensure sufficient time to help participants who needed to apply for visas, and/or whose travel we needed to support. Indeed, visa applications can take a significant amount of time. For example, one participant had a visa issued only two days before travel. For similar reasons, it was necessary to book travel for fully-funded participants before visa applications were completed, otherwise flight costs could have risen to unaffordable rates as the travel dates approached.
@@
The only costs that fully-funded participants were required to pay for themselves up-front, and later obtain a reimbursement, were the visa costs, travel insurance (as often not provided by the home universities), and local travel costs. We note that costs associated with obtaining visas were sometimes quite high, due to participants needing to travel to consulates (sometimes multiple times). Although this process generally went well, some applicants had no access to personal bank accounts and arrangement had to be made with a third party to make the reimbursement because cash payments were not possible. In some cases, we were not aware of this until after it was too late to help with alternative methods of payment beforehand.
@@
The participants applied for, and received, standard visitor visas. These are valid for six months, and some participants organised other work visits while they were in the UK. This required some extra coordination, but we were happy to support it. We had some requests during the meeting, or after, for other work visits and it would have been useful to highlight this opportunity to participants beforehand. This could have allowed participants to make the best use of their visas while they were valid. However, at the time, we caution that while visits to the UK for conferences and meeting did not require approval from the Academic Technology Approval Scheme (ATAS), research visits of any length did require approval. If people wish to join a visit to carry out research along with this, we recommend they apply for ATAS certification, as required, at the same time or even before the visa.
@@
The following sub-sections describe the activities and their impact, as much as we are able to assess at the time of the workshop.
@@
Around one month before the workshop, we created a collaborative online Slack space, to allow participants to start introducing themselves to each other. Additionally, to encourage participants getting to know each other before the conference, we asked participants to create two minute introductory videos. These videos were shared on a private YouTube playlist in the month leading up to the conference. Participants were provided with clear guidelines on how to record these videos on a phone, and what these videos should include (i.e., their name, affiliation, research interests, and what they hope to get out of the conference). Only 35/107 participants created these videos; with mostly the more junior participants participating (the LOC and SOC were all compelled to make a video). Nonetheless, we had several positive comments during the feedback questionnaire about the introductory videos being a good idea, and useful to start to get to know fellow participants.
@@
We split into groups of $\approx$25, across separate rooms. These groups were then re-shuffled half way through the session. Before the activity all participants had been assigned their groups/rooms so that they knew where to be. Two activities were used: (1) “Find 10 things in common” and (2) “Queen Bee: Find the Answer to my Questions”. The former activity had each group split into 2-3 sub-groups and then take turns to briefly introduce themselves to each other. They then had ten minutes to try and find ten things in common across the whole group. This could be related to research tools used, hobbies, favourite types of food etc. The second activity involved sub-groups of $\approx$3–5. One person in each group is nominated the “Queen Bee”, who gives their “worker bees” a question that they want the answer to. This could be something scientific (e.g., “who could I talk to about processing these type of astronomical data”), or anything else (e.g., “what’s the best restaurant to get good Indian food in Newcastle”). All the worker bees, from each group, then spend roughly three minutes speaking to everyone else in the room to get the answers, and report back to the Queen Bee. The Queen Bees could then report back to the whole room what they found out, before somebody else takes a turn as Queen Bee.
The top-left panel of Figure 2 shows the results of asking participants “How did you find the networking session on Monday?”, with answers between 1 for poor and 5 for excellent. It can be seen that the activity was extremely popular across all participants, with 91% scoring 4 or 5. In Figure 2 we also split the responses for only the African affiliated/origin participants, and found equal levels of positive responses for this specific group. The qualitative comments suggested the activity was very effective for helping early-career researchers feel more relaxed and able to approach people for conversations throughout the week. It was especially noticeable feeling more comfortable approaching senior academics after this activity. Some other comments suggested the activity was a bit tiring, so it could have been reduced in length and/or more rest time built in to the programme after this intense activity.
@@
Sparkler talks were 90 second presentations (with one slide only). These were done over two scheduled blocks on the first two days. The purpose of these were to allow participants to introduce themselves, their main research interest and/or one key scientific result, and to advertise a poster (if relevant). As described above, participants were given examples and training on how to effectively put these together, with a focus on avoiding participants trying to give a full scientific talk very quickly!
The SOC selected chairs for the contribution sessions, to be more senior academics, but keeping the geographical representation the same as the overall conference. We provided the chairs with moderator guidelines, which most chairs adhered to. These included information about keeping to time, encouraging a diversity of participants to ask questions, and ensuring the code of conduct is upheld. The LOC provided technical support for each session, with specific LOC members identified for each session.
@@
The feedback we received indicates that we achieved our objective of creating a collaborative environment where people “across continents” could come together to network and share in scientific discussion. The bottom-right panel of Figure 2 shows the results of the participant feedback to the question: “How comfortable did you feel to have discussions/conversations with new people during the conference (e.g., during coffee breaks, lunchtimes…)”, ranging from 5 (Very) to 1 (Not at all). The results are overall positive with 55% scoring 5 and 36% scoring 4, with a similar positive response for only the African affiliated/origin groups (62% scoring 5 and 36% scoring 4).
@@
We received very little feedback from the small number of online participants. However, we know that there they mostly only participated during the scientific contributions part of the conference. Our focus was always on generating an effective in-person conference experience. Future consideration would be needed to make a more effective hybrid conference format to facilitate networking, and/or to consider an online-only equivalent.
@@
Immediately after the conference, initial signs were already encouraging for a longer-term impact of the conference. For example, when we asked “Would you be interested in attending a conference like this in a future?”, (see Figure 3), 94% of all participants answered “yes” (100% of the African affiliated/origin participants). There was enthusiasm to run another similar meeting, potentially on the continent of Africa next time. Indeed, there are now provisional plans for an AGN themed conference to occur in Kenya in September 2026.
One year on: longer-term impact of the conference
One year after the AGN Across Continents conference (i.e., in July 2025), we contacted the participants to ask about longer-term impacts that they had experienced due to the conference. The responses came in two forms: (1) non-anonymous informal email descriptions of impacts and (2) anonymous answers to a structured form (completed by 19 participants). The latter included two main parts; firstly asking the respondents to state which of the listed impacts they had experienced (see Table 1) and secondly, open-ended requests for benefits that experienced on either research projects or their professional development.
Our analysis finds clear evidence of long-term benefits in research collaboration, professional development, and community-building, especially for African-affiliated early-career participants. Quantitatively, a majority of respondents reported initiating new projects or ideas, while nearly all reported concrete career gains (see Table 1). Qualitative feedback highlights multiple joint research projects launched (e.g., a new AGN program at the Ghana observatory), mentorship connections formed, and increased confidence and visibility among early-career researchers. Many participants (including one who secured a PhD supervisor from conference contacts) attribute career boosts to the conference. The “Buddy Scheme” introduced cross-continental mentoring, with many buddies remaining in contact, reflecting sustained networking. Feedback emphasises that greater African representation made senior researchers more approachable, benefitting students and early-career scientists alike. These outcomes align with the conference’s inclusivity goals, by “bringing researchers together from Africa, Europe, and beyond”, the meeting fostered a more global approach to science and inclusivity.
Key metrics from the participant survey show high levels of follow-up engagement and career benefits:
New Research Collaborations: 58% of respondents reported discussing project ideas with the intent to start a collaboration; 16% have already started a specific joint project; 16% discussed ideas but “nothing came of them” yet; and only 10% reported no new project discussions.
Professional Development: All respondents (100%) indicated that the conference gave them new research ideas, greater readiness to attend future conferences, new professional contacts, and that they listed the conference on their CV [^7] The authors anecdotally note we have indeed seen this listed on CVs in our roles of assessing various applications.. 74% learned of new job or funding opportunities through the conference, and 37% have already applied for positions they discovered through these contacts. These trends are summarized in Table 1.
Table 1:
Research Collaboration Outcomes Total (19) %
Discussed project ideas (hoping to start) 11 58%
Started a specific joint project 3 16%
Discussed ideas (no project started) 3 16%
No project discussions 2 10%
Professional Development Outcomes Total (19) %
New ideas to advance future research 19 100
Learned of new job/funding opportunities 14 74%
Applied for jobs (through contacts/opportunities) 7 37%
New contacts/networking established 19 100%
More prepared to attend future conferences 16 84%
Conference attendance on CV 19 100%
Case studies of some of the outcomes
1. A multi-country collaboration: Pan-African AGN survey using SALT/IFS (PanAfroAGN-SI): Shortly after the conference, an international team including African (mostly) and European astronomers proposed a major SALT observing programme on AGN. This project built on contacts made at the workshop, and it will be led by African scientists. The programme embodies the conference theme, covering key AGN science questions, linking African telescope resources with broad expertise [7]. The project has been approved as a Large Program (3–4 years of observations), with a provisional 10 hours as a pilot conducted initially, with 100s of hours expected to follow. Data reduction of the first observations is underway.
2. UKRI-funded GGOS-Africa project: During the conference , the project co-ordinator (Jack Radcliffe) met future project partners for the Africa Global Geodetic Observing System (GCOS) application to build geodetic stations across Africa. The proposal was successful, funded by UKRI (with grant code UKRI476) and the project is ongoing. [^8] https://www.jb.man.ac.uk/GGOS-Africa/overview.html
3. East African AGN Paper: A graduate student from Rwanda combined efforts with peers in Uganda and Ethiopia to co-author a manuscript on the stellar populations of AGN. The idea emerged from networking at the conference and a follow-up virtual meeting. This new collaboration illustrates the shared expertise and data resources that participants foresaw during the post-conference discussions.
4. DARA-STFC Postdoc fellowship placement: A PhD student from Africa attributes their current postdoc position as a result of their “collaborative discussions at the conference”.
5. Strengthening African Research Capacity via “Buddy Collaboration”: A clear career-development success emerged from the Early-Career Researcher “Buddy Scheme”. Through the buddy pairing, Ghanaian researcher Dr. Proven Emmanuel established a collaboration with Dr. Evaristus Iyida, resulting in a DARA postdoctoral appointment at the Ghana Radio Astronomy Observatory, significantly contributing to the institute’s scientific output and its capacity for AGN-focused research. This case highlights how intentional mentorship and cross-continental networking directly support institutional growth and long-term skills development in the African astronomy community.
6. Research visit to Ghana by a senior European-based professor: following discussions during the conference, a senior professor, Leah Morabito, was invited to deliver DARA lectures at the Ghana Radio Astronomy observatory. The local researchers at the observatory note that the involvement of this professor has improved the quality and the delivery of these astronomy lectures significantly.
7. Durham visit (mentored research experience): Another early-career researcher arranged a three month research visit to Durham University, UK, to learn radio techniques. They attribute this to an introduction made by a conference presenter. The visit included attending a regional workshop in South Africa, strengthening "within Africa" ties. The participant reported that discussing a potential research visit at the conference put them in contact with a supportive mentor, directly benefiting their PhD research trajectory. A similar arrangement is being made for another participant to visit Durham University in early 2027.
8. Support of masters/undergraduate transitioning to higher degrees The authors have a couple of examples of how junior conference participants have been supported in their transition to higher degrees, thanks to the connections made at the conference. For example, one of the African conference participants is now studying at Durham for an Master of Science by Research, supported via DARA. Another African student was successful in obtaining a PhD in Cork after support and mentorship in both searching for opportunities and preparing the application materials. The meeting led to me establishing a new collaborative project using SDSS-V data between Dr James Aird (Edinburgh University) and Dr Eli Kasai (University of Namibia), which then led to the recruitment of a Ghanian student (Mohammed Iddrisu Nlowie) through the DARA project to complete an MSc by Research in Edinburgh.
@@
Other smaller examples include two PhD students (one of African origin) deciding to initiate a new project combining the data of one student with the methods of the other student, and a student visit to Durham to start a radio project (combined with a workshop). It is likely other examples occurred but we do not have complete information.
This series of examples highlights the long-term impact such a conference can have on individuals and communities. It is still too early to tell the long-term scientific impacts of this conference, as the projects/PhD programmes etc. are only just getting started. We did not take a very systematic approach to tracking long-term impacts, so there may be more examples we have missed here. Furthermore, it is clear that mentorship from new networks/relationships is a key factor in the benefits of this workshop. Therefore, on reflection we could have done more to help sustain the “Buddy Scheme” and potentially other mentorship relationships after the conference.
Key recommendation: build in a methodology to sustain and track impacts from the workshop on a long-term time scale and consider funding possibilities for seeding ideas developed. Consider a way to sustain longer term “Buddy Scheme” or new mentorship relationships after the workshop.
@@
The main goal of this conference was to hold a scientific meeting, discussing timely scientific problems, but with a broader impact to integrate African researchers into the international community and to strengthen ties between the European and African scientific communities. According to the feedback from one participant, they reached their scientific goals for the conference, along with the bonuses of “1. unplanned research insights and collaboration directions …; 2. appreciating the responsibility that comes with the privilege of being West-affiliated; 3. getting to know African astronomers who are great role models for everyone.” From a different perspective, “Having greater representation of Africans really made people much more approachable.” The social activities allowed everyone to bond, while many people noted that “The quality of the participants’ science contributions has been really impressive.” The feedback has been overwhelmingly positive, and there is appetite for another conference in the community. We seem to have struck the right balance between science and social aspects about right: as one participant noted the conference was “really well thought out in terms of inclusivity, time table, networking opportunities, making people (especially students) feel comfortable to attend and talk to more senior researchers, super positive and supportive atmosphere, extremely interesting and high-quality science.”
@@
We have learned a lot from this entire process, from simple things like: what people actually have to pay for if applying for a visa and the logistical considerations we take for granted that may be different in other countries. Even the convention of putting a first name before, or after, a family name differs across countries. The diversity of the LOC and SOC were fundamental in identifying potential issues before they arose, and problem-solving when unexpected challenges came up. The biggest challenges revolved around ensuring our budget would stretch to cover all of the costs we promised, and our key recommendation here is that funding requests start early, and that as much as possible should be paid for directly by the host organisations.
@@
The funding for this conference pulled together from several sources. It started from the fact that the co-chairs had small amounts from their fellowships for modest science meetings, which could be combined for a single, more impactful meeting. Other funding was necessary, and it is crucial to start looking for funding sources early ($\approx$18–24 months before).
@@
Overall, this meeting exceeded our expectations. The LOC and SOC did an incredible job to enact the vision, and to ensure that every small aspect was considered and addressed. Before and during the conference, the engagement from all participants was absolutely amazing. It was wonderful to see senior and junior researchers from Africa, Europe and beyond, interacting and discussing science. We hope that the collaborations that have started will bring our communities together for a more global approach to science and inclusivity.
Reflecting on the impacts we have managed to document a year after the conference, we conclude that the AGN Across Continents conference has had a measurable, lasting impact. It seeded new research collaborations, advanced careers, and strengthened a diverse, inclusive community of astronomers. The overwhelmingly positive feedback and ongoing collaborations demonstrate that the event not only met its immediate objectives but also laid foundations for sustainable African–European research network. Future initiatives should build on these successes by supporting follow-up collaborations, expanding mentoring schemes, and continuing to prioritize diversity and inclusion in academic conferences.